OpenStudy (anonymous):

f(x) = Square root of quantity 3x+7 g(x) = Square root of quantity 3x-7 Find (f + g)(x).

OpenStudy (anonymous):

I messed up on my first answer so let me try again

OpenStudy (anonymous):

haha ok thanks

OpenStudy (anonymous):

x= -7/3 if f(x)=square root of 3x+7 x=7/3 if g(x)= square root of 3x-7

OpenStudy (anonymous):

So, I think you would just plug in x for each equation, and plug those numbers in for (f+g)(x)

OpenStudy (anonymous):

if these are my options: a)square root of six x. b) x Square root of six. c) 3x d)Square root of quantity three x plus seven. +Square root of quantity three minus plus seven. would it be a?

OpenStudy (anonymous):

@blueberry27

OpenStudy (anonymous):

The answer choices look a bit confusing. Why do you think it's a?

OpenStudy (anonymous):